In a study, the sys5olic blood pressure for population of female participants are found to have a mean of 120 and a standard deviation of 12. Suppose the the blood pressure of a random sample of 64 female participants is obtained from this study. What is the probability that the sample mean is between 118 and 123?
Answer by stanbon(75887)

In a study, the systolic blood pressure for population of female participants are found to have a mean of 120 and a standard deviation of 12. Suppose the the blood pressure of a random sample of 64 female participants is obtained from this study. What is the probability that the sample mean is between 118 and 123?
-------
z(118) = (118-120)/[12/sqrt(64)] = -4/3
----
z(123) = (123-120)/[12/sqrt(64)] = 2
----
P(118< x-bar < 123) = P(-4/3 < z < 2) = 0.8860
===============================
Cheers,
Stan H.